The Tinsley Company exchanged land that it had been holding for future plant expansion for a more suitable parcel located farther from residential areas. Tinsley carried the land at its original cost of $87,500. According to an independent appraisal, the land currently is worth $210,000. Tinsley paid $30,000 in cash to complete the transaction.

Prepare the journal entry to record the exchange except that Tinsley received $42,000 in the exchange, and the exchange lacks commercial substance.
Answer #1

SOLUTION

Fair value of new land = $210,000

Carrying value of old land = $87,500

Realized gain = 210,000 - 87,500 = 122,500

Cash received = $42,000

Here, percentage of cash to total consideration = 42,000 / 210,000 = 20%

Recognized gain = 20% * 122,500 = 24,500

Journal entry-

S.No. Accounts titles and Explanation Debit ($) Credit ($)
1. Land- New 70,000
Cash 42,000
Land- Old 87,500
Gain 24,500
(Being old Land exchanged for new Land and recognizing gain in the transaction)